Okay, so I'm working on Aleks, and they all of the sudden switched the problems I was working on and didn't give me an example on how to solve questions like these. I'm completely lost and stuck while trying to figure this out. Could someone please explain to me how I solve these types of questions?
Two factory plants are making TV panels. Yesterday, Plant A produced twice as many panels as Plant B. Four percent of the panels from Plant A and 3%of the panels from Plant B were defective. How many panels did Plant B produce, if the two plants together produced 660 defective panels?

Answer by mananth(16946)


Let B produce x panels
Then A produces 2x panels
Four percent of the panels from Plant A and 3%of the panels from Plant B were defective.
0.04(2x) +0.03x = 660
0.08x +0.03x = 660
0.11x = 660
/0.11
x = 660/0.11
x= 6600
Panels produced by B are 6600
CHECK
4%(6600)+3%(13200) =660